PHYSICS
1. A piece of copper block of mass 24g at 2300C is placed in a copper calorimeter of
mass 60g containing 54g of water at 31oC. Assuming heat losses are negligible,
calculate the final steady temperature of the mixture. [Specific hest capacity of water
= 4200Jkg – 1 K – 1
, specific heat capacity of copper = 400Jkg – 1 K – 1
].
2. The lower and upper fixed points marked on a mercury in glass thermometer are
210mm apart. The end of the mercury column in the tube is 49mm above the lower
fixed point in a room. What is the temperature of the room in degree celcius.
1
Expert's answer
2020-04-02T10:50:14-0400

1)

Heat lost by the copper block = heat gained by water + heat gained by calorimeter

Let the equilibrium temperature be T

"0.024\times400\times(210-T)=0.06\times400\times(T-31)+0.054\times4200\times(T-31)"

"2016-9.6T=24T-744+226.8T-7030.8"

"260.4T=5758.8
\\T=22.12^oC"


2)

"210mm .... 100^oC\\
49mm .... (\theta)\\

0mm .... 0^oC\\

\frac{49 - 0}{210 - 0} =\frac{\theta - 0}{100 - 0}

\\\frac{49}{210} = \frac{\theta}{100}
\\
\theta = 23.3oC"